HIGH-RESOLUTION INFRARED SPECTROSCOPY OF HCN-Zn_{n} CLUSTERS

\underline{P. L. STILES}, and R. E. MILLER, Department of Chemistry, The University of North Carolina at Chapel Hill, Chapel Hill, NC 27599.

Rotationally resolved infrared spectra have been obtained for the first time for tr ansition metal complex-adsorbate systems. The zinc clusters are formed in helium d roplets, to which are added a single HCN molecule. The C-H stretches of the comple xes have been measured for HCN-Zn_{n}, where n=1-4. The rotaional constants and di pole moments are reported and compared to ab initio calculations.